Bob can paint a small house in 8 hours. Joe could paint the same size house in 10 hours. On Friday, Bob started painting the house by himself. he had to leave after 2 hours and was replaced by Joe. How long did it take joe to finish painting?
Answer by hamsanash1981@gmail.com(151)

Bob's 1 hour work = 1/8
Joe's 1 hour work = 1/10
Bob's 2 hours work = 2/8 = 1/4
remaining part of work for 1 hour = 1-1/4 = 3/4
Time to finish work is 4/3 hours.
therefore work done by Joe 4/3 = 4/3 * 1/10
= 2/15
Joe took = 15/2 hours or 7 hours 30 mins to complete the work.